# Linearization of ergodic McKean SDEs and applications

### v1

@@ 3.1 Maximum likelihood estimation
We consider the case where the confining and interaction potentials $V$ and $W$ depend on unknown parameters. Our goal is to infer these parameters by observing a trajectory of the mean field SDE (1.2). Hence, consider the equation on the whole space $\mathbb{R}^{d}$ and in the time horizon $[0,T]$
@@
We now define the linearized version of the log-likelihood around the invariant measure $f_{\infty}$,
@@ 3.1.1 Asymptotic unbiasedness of the linearized MLE
    $\theta\mapsto\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x;\theta)\,\mathrm{d}x$ (3.12)
@@
The condition in Assumption 3.1(iii) is necessary to guarantee the identifiability of the unknown parameter $\theta_{0}$, and it is related to the nondegeneracy of the Fisher information matrix in [24].
@@
Under Assumptions 2.3 and 3.1, it holds
@@
Our theoretical analysis is inspired by the proof of [11, Corollary 4.1], where the linear case is treated. In fact, in our approach the likelihood function is computed from the linear process, and the nonlinearity only appears in the observation model. The proof of Theorem 3.3 relies on the following preliminary results.
@@
Under Assumptions 2.3 and 3.1, it holds for all $\theta\in\Theta$
@@
Under Assumptions 2.3 and 3.1, it holds
@@
The desired result follows from Lemma 3.5 by replacing $\theta$ with $\theta_{0}$. ∎
@@
Under Assumptions 2.3 and 3.1, it holds
@@
which is uniformly continuous in $\theta$ on $\Theta$, uniformly in $T\in[1,\infty)$, with probability $1$, by [11, Theorem 3.1]. Then, by Lemma 3.5 and the strong law of large numbers for martingales we deduce that for all $\theta\in\Theta$
@@
Under Assumptions 2.3 and 3.1, there exists a constant $L>0$, independent of $T$, such that for all $\theta_{1},\theta_{2}\in\Theta$
@@
where the uniform bound in time of the last inequality is due to Lemma 3.5. By the definition of $\widetilde{b}_{\theta}$ in equation (3.4) we obtain
@@
Moreover, since $\nabla V$ and $\nabla W$ are Lipschitz in $\theta$ by Assumption 3.1(ii) and polynomially bounded in $x$, using the uniform boundedness of the moments and the ergodic theorem, and due to the differentiability of the invariant measure with respect to $\theta$ [9], we have that there exists a constant $L>0$, independent of $T$, such that
@@
We can finally prove Theorem 3.3.
Proof of Theorem 3.3.
@@
Applying Corollaries 3.6 and 3.7 we obtain
@@
Let us now define the function $h=h(T,\theta)$ as in equation (3.25), which is Lipschitz in $\theta$, uniformly in $T$, by Lemma 3.8, and notice that
@@
which by definition (3.25) and Lemma 3.5 coincides with
$\Omega=\left\{\theta\in\Theta\colon\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x)\,\mathrm{d}x=0\right\}.$ (3.36)
Finally, by the identifiability condition in Assumption 3.1(iii) we get that $\Omega=\{\theta_{0}\}$, and the desired result follows. ∎
@@ 3.1.2 Numerical examples
Figure 1: Comparison between the MLE $\widehat{\theta}_{T}$ and the linearized MLE $\widetilde{\theta}_{T}$ for different values of the final time of observation $T\in[0,10^{3}]$, considering the quadratic interaction potential and two different examples of confining potentials.
@@
In the first plot in Figure 1, we compare the two estimators for different values of the final time of observation $T\in[0,10^{3}]$, and notice that, except for a short time interval, the difference between $\widehat{\theta}_{T}$ and $\widetilde{\theta}_{T}$ is negligible. In particular, they both provide an accurate approximation of the exact unknown.
@@
$\mathrm{d}Y_{t}$ $=-\theta X_{t}^{3}\,\mathrm{d}t-(1-\theta)X_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$
@@
which does not need the expectation $\operatorname{\mathbb{E}}[X_{t}]$. We note that, in this case, McKean’s SDE must be approximated through the interacting particle system (1.1). In particular, we use $N=500$ particles, and a path of the mean field SDE is approximated by a path of one of the trajectories of the particle system. Due to exchangeability, it does not matter which particle we choose. We notice that assuming to know the expectation $\operatorname{\mathbb{E}}[X_{t}]$ at all times implies that all the interacting particle system must be observed. On the contrary, for the linearized MLE, only the trajectory of a single particle is required. The behavior of the two estimators for different values of the final time of observation $T\in[0,10^{3}]$ is shown in the second plot in Figure 1. As expected, the two estimators provide us with almost identical values for the estimated parameter, after a short initial transient time.
@@ 3.2 The diffusive-mean field limit
We revisit the problem of studying the joint diffusive-mean field limit for weakly interacting diffusions in periodic confining and interaction potentials that was studied in [22]. As discussed in that paper, the two limits of sending $\varepsilon\to 0$ and $N\to\infty$ commute only when the mean field dynamics on the torus does not exhibit phase transitions. Here, we study only the “easy” order of taking the two limits, first $N\to\infty$ and then sending $\varepsilon$ to 00, and only under the assumption that the mean field SDE has a unique steady state. A full analysis of the problem based on the linearized SDE, for both the overdamped and the underdamped dynamics, will be presented elsewhere.
@@ 3.2.1 Limit distribution of the linearized process
Let $Y_{t}$ be the solution of the SDE (3.58), and let $\widetilde{Y}^{\varepsilon}_{t}=\varepsilon Y^{\varepsilon}_{t/\varepsilon^{2}}$. Then, it holds
@@
Reasoning as in the proof of Theorem 3.9, we obtain
@@ 3.2.2 Numerical example
Figure 2: Numerical verification of the CLTs in equation (3.57) and Theorem 3.9, considering sinusoidal confining and interaction potentials.
We numerically verify in a one-dimensional test case the CLTs for the nonlinear and linear processes $X_{t}$ and $Y_{t}$ given in equations (3.57) and Theorem 3.9, respectively. We consider the setting of [22, Lemma 1.26] below the phase transition, i.e., we set $W(x)=-\cos(2\pi x)$, $V(x)=-\xi\cos(2\pi x)$ with $\xi=1/2$, and $\beta=1$. In [22, Lemma 1.26], it is proved that the wrapping of the solution of the McKean SDE has a unique invariant measure with density
@@
In Figure 2 we show the histograms of $X_{t}/\sqrt{t}$ and $Y_{t}/\sqrt{t}$ with $t=10^{3}$ for $500$ different realizations of the Brownian motion. We remark that, differently from equation (3.76) that can be simulated, the McKean SDE for $X_{t}$ must be approximated by the interacting particle system (1.1), for which we use $N=250$ particles. All the SDEs have initial distribution $\mu_{0}=\nu_{0}=\delta_{0}$, and they are discretized using the Euler–Maruyama method with a time step $\Delta t=10^{-2}$. We observe that the histograms match the limit Gaussian distribution $\mathcal{N}(0,2\beta^{-1}\mathcal{D})$ predicted by the theory.

### v2

@@ 3.1 Maximum likelihood estimation
[…]
@@
We now define the linearized version of the log-likelihood around the invariant measure $f_{\infty}(\cdot;\theta)$,
@@ 3.1.1 Asymptotic unbiasedness of the linearized MLE
    $\theta\mapsto\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x;\theta_{0})\,\mathrm{d}x$ (3.12)
@@
The condition in Assumption˜3.1(iii), which appears also in [63], is necessary to guarantee the identifiability of the unknown parameter $\theta_{0}$, and it is related to the nondegeneracy of the Fisher information matrix in [24]. We notice that this condition is satisfied, e.g., if $\widetilde{b}_{\theta}$ is an affine function of $\theta$.
@@
Under Assumptions˜2.3 and 3.1, it holds
@@
Our theoretical analysis is inspired by the proof of [11, Corollary 4.1], where the linear case is treated. In fact, in our approach the likelihood function is computed from the linear process, and the nonlinearity only appears in the observation model. The proof of Theorem˜3.3 relies on the following preliminary results.
@@
Under Assumptions˜2.3 and 3.1, it holds for all $\theta\in\Theta$
@@
Under Assumptions˜2.3 and 3.1, it holds
@@
The desired result follows from Lemma˜3.5 by replacing $\theta$ with $\theta_{0}$. ∎
@@
Under Assumptions˜2.3 and 3.1, it holds
@@
which is uniformly continuous in $\theta$ on $\Theta$, uniformly in $T\in[1,\infty)$, with probability 111, by [11, Theorem 3.1]. Then, by Lemma˜3.5 and the strong law of large numbers for martingales we deduce that for all $\theta\in\Theta$
@@
Under Assumptions˜2.3 and 3.1, there exists a constant $L>0$, independent of $T$, such that for all $\theta_{1},\theta_{2}\in\Theta$
@@
where the uniform bound in time of the last inequality is due to Lemma˜3.5. By the definition of $\widetilde{b}_{\theta}$ in equation (3.4) we obtain
@@
Moreover, since $\nabla V$ and $\nabla W$ are Lipschitz in $\theta$ by Assumption˜3.1(ii) and polynomially bounded in $x$, using the uniform boundedness of the moments and the ergodic theorem, and due to the differentiability of the invariant measure with respect to $\theta$ [9], we have that there exists a constant $L>0$, independent of $T$, such that
@@
We can finally prove Theorem˜3.3.
Proof of Theorem˜3.3.
@@
Applying Corollaries˜3.6 and 3.7 we obtain
@@
Let us now define the function $h=h(T,\theta)$ as in equation (3.25), which is Lipschitz in $\theta$, uniformly in $T$, by Lemma˜3.8, and notice that
@@
which by definition (3.25) and Lemma˜3.5 coincides with
$\Omega=\left\{\theta\in\Theta\colon\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x;\theta_{0})\,\mathrm{d}x=0\right\}.$ (3.36)
Finally, by the identifiability condition in Assumption˜3.1(iii) we get that $\Omega=\{\theta_{0}\}$, and the desired result follows. ∎
@@ 3.1.2 Numerical examples
Figure 1: Comparison between the MLE $\widehat{\theta}_{T}$ and the linearized MLE $\widetilde{\theta}_{T}$ for different values of the final time of observation $T\in[0,10^{3}]$, considering the quadratic interaction potential and two different examples of confining potentials, when the unknown coefficient is in the confining potential.
Figure 2: Comparison between the MLE $\widehat{\theta}_{T}$ and the linearized MLE $\widetilde{\theta}_{T}$ for different values of the final time of observation $T\in[0,10^{3}]$, considering the quadratic interaction potential and two different examples of confining potentials, when the unknown coefficient is in the interaction potential.
@@
In the first plot in Figure˜1, we compare the two estimators for different values of the final time of observation $T\in[0,10^{3}]$, and notice that, except for a short time interval, the difference between $\widehat{\theta}_{T}$ and $\widetilde{\theta}_{T}$ is negligible. In particular, they both provide an accurate approximation of the exact unknown.
@@
$\mathrm{d}Y_{t}$ $=-\theta Y_{t}^{3}\,\mathrm{d}t-(1-\theta)Y_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$
@@
which does not need the expectation $\operatorname{\mathbb{E}}[X_{t}]$. We note that, in this case, McKean’s SDE must be approximated through the interacting particle system (1.1). In particular, we use $N=500$ particles, and a path of the mean field SDE is approximated by a path of one of the trajectories of the particle system. Due to exchangeability, it does not matter which particle we choose. We notice that assuming to know the expectation $\operatorname{\mathbb{E}}[X_{t}]$ at all times implies that all the interacting particle system must be observed. On the contrary, for the linearized MLE, only the trajectory of a single particle is required. The behavior of the two estimators for different values of the final time of observation $T\in[0,10^{3}]$ is shown in the second plot in Figure˜1. As expected, the two estimators provide us with almost identical values for the estimated parameter, after a short initial transient time.
We finally remark that similar results are obtained if the unknown parameter to be estimated appears in the interaction potential. Specifically, consider the mean field Ornstein–Uhlenbeck process together with its linearized process
$\mathrm{d}X_{t}$ $=-X_{t}\,\mathrm{d}t-\theta(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.48)
$\mathrm{d}Y_{t}$ $=-(\theta+1)Y_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$
In this case we have $\operatorname{\mathbb{E}}[X_{t}]=\operatorname{\mathbb{E}}[X_{0}]e^{-t}$ and the MLE estimators are
$\widehat{\theta}_{T}$ $=-\frac{\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{0}]e^{-t})\,\mathrm{d}X_{t}+\int_{0}^{T}X_{t}(X_{t}-\operatorname{\mathbb{E}}[X_{0}]e^{-t})\,\mathrm{d}t}{\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{0}]e^{-t})^{2}\,\mathrm{d}t},$ (3.49)
$\widetilde{\theta}_{T}$ $=-\frac{\int_{0}^{T}X_{t}\,\mathrm{d}X_{t}}{\int_{0}^{T}X_{t}^{2}\,\mathrm{d}t}-1.$
On the other hand, regarding the bistable confining potential example, the linear and nonlinear processes satisfy
$\mathrm{d}X_{t}$ $=-(X_{t}^{3}-X_{t})\,\mathrm{d}t-\theta(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.50)
$\mathrm{d}Y_{t}$ $=-Y_{t}^{3}\,\mathrm{d}t-(\theta-1)Y_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$
and the corresponding MLE estimators read
$\widehat{\theta}_{T}$ $=-\frac{\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}X_{t}+\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{t}])(X_{t}^{3}-X_{t})\,\mathrm{d}t}{\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{t}])^{2}\,\mathrm{d}t},$ (3.51)
$\widetilde{\theta}_{T}$ $=-\frac{\int_{0}^{T}X_{t}\,\mathrm{d}X_{t}+\int_{0}^{T}X_{t}(X_{t}^{3}-X_{t})\,\mathrm{d}t}{\int_{0}^{T}X_{t}^{2}\,\mathrm{d}t}.$
The numerical results are reported in Figure˜2, where we observe that the linear and nonlinear MLE estimators agree and tend to coincide when the final time of observation increases.
@@ 3.2 The diffusive-mean field limit
We revisit the problem of studying the joint diffusive-mean field limit for weakly interacting diffusions in periodic confining and interaction potentials that was studied in [22]. As discussed in that paper, the two limits of sending $\varepsilon\to 0$ and $N\to\infty$ commute only when the mean field dynamics on the torus does not exhibit phase transitions. Here, we study only the “easy” order of taking the two limits, first $N\to\infty$ and then sending $\varepsilon$ to 0, and only under the assumption that the mean field SDE has a unique steady state. A full analysis of the problem based on the linearized SDE, for both the overdamped and the underdamped dynamics, will be presented elsewhere.
@@ 3.2.1 Limit distribution of the linearized process
Let $Y_{t}$ be the solution of the SDE (3.62), and let $\widetilde{Y}^{\varepsilon}_{t}=\varepsilon Y_{t/\varepsilon^{2}}$. Then, it holds
@@
Reasoning as in the proof of Theorem˜3.9, we obtain
@@ 3.2.2 Numerical example
Figure 3: Numerical verification of the CLTs in equation (3.61) and Theorem˜3.9, considering sinusoidal confining and interaction potentials.
We numerically verify in a one-dimensional test case the CLTs for the nonlinear and linear processes $X_{t}$ and $Y_{t}$ given in equations (3.61) and Theorem˜3.9, respectively. We consider the setting of [22, Lemma 1.26] below the phase transition, i.e., we set $W(x)=-\cos(2\pi x)$, $V(x)=-\xi\cos(2\pi x)$ with $\xi=1/2$, and $\beta=1$. In [22, Lemma 1.26], it is proved that the wrapping of the solution of the McKean SDE has a unique invariant measure with density
@@
In Figure˜3 we show the histograms of $X_{t}/\sqrt{t}$ and $Y_{t}/\sqrt{t}$ with $t=10^{3}$ for 500500500 different realizations of the Brownian motion. We remark that, differently from equation (3.80) that can be simulated, the McKean SDE for $X_{t}$ must be approximated by the interacting particle system (1.1), for which we use $N=250$ particles. All the SDEs have initial distribution $\mu_{0}=\nu_{0}=\delta_{0}$, and they are discretized using the Euler–Maruyama method with a time step $\Delta t=10^{-2}$. We observe that the histograms match the limit Gaussian distribution $\mathcal{N}(0,2\beta^{-1}\mathcal{D})$ predicted by the theory.